# AI-Driven Early Mental Health Screening with Limited Data: Analyzing Selfies of Pregnant Women

### v1

@@ Abstract
Major Depressive Disorder and anxiety disorders affect millions globally, contributing significantly to the burden of mental health issues. Early screening is crucial for effective intervention, as timely identification of mental health issues can significantly improve treatment outcomes. Artificial intelligence (AI) can be valuable for improving the screening of mental disorders, enabling early intervention and better treatment outcomes. AI-driven screening can leverage the analysis of multiple data sources, including facial features in digital images. However, existing methods often rely on controlled environments or specialized equipment, limiting their broad applicability. This study explores the potential of AI models for ubiquitous depression-anxiety screening given face-centric selfies. The investigation focuses on high-risk pregnant patients, a population that is particularly vulnerable to mental health issues. To cope with limited training data resulting from our clinical setup, pre-trained models were utilized in two different approaches: fine-tuning convolutional neural networks (CNNs) originally designed for facial expression recognition and employing vision-language models (VLMs) for zero-shot analysis of facial expressions. Experimental results indicate that the proposed VLM-based method significantly outperforms CNNs, achieving an accuracy of 77.6% and an F1-score of 56.0%. Although there is significant room for improvement, the results suggest that VLMs can be a promising approach for mental health screening, especially in scenarios with limited data.
@@ 1 Introduction
Major depressive disorder (depression) is a prevalent mental health disorder affecting over 280 million people globally and a leading cause of disability worldwide. It contributes significantly to the global disease burden [1]. Similarly, anxiety disorders, marked by excessive worry and fear, impact millions of lives, often leading to debilitating effects on daily functioning and quality of life. Particularly, there is a recognized vulnerability to mood disorders during pregnancy, with possible worsening of previous emotional conditions and the emergence of new altered mental states that increase the risk of depression and anxiety [2]. This vulnerability is certainly influenced by the significant hormonal changes of pregnancy. However, it is also important to recognize that this is a time of great physical, psycho-emotional, cultural, and social changes, generating psychic stress and increased anxiety.
Several studies have pointed to the association between depression and anxiety during pregnancy with unfavorable obstetric and neonatal outcomes. These disorders can increase the risk of obstetric complications such as cesarean delivery, preeclampsia, preterm birth, low birth weight, small for gestational age newborns, and newborns with low Apgar scores, indicating lower oxygenation at birth [3, 4, 5, 6]. These findings underline the importance of making a correct and early diagnosis, allowing for appropriate psychiatric and psychological follow-up during pregnancy to improve both maternal and neonatal outcomes.
For early intervention and improved treatment outcomes, screening processes are essential for identifying individuals who may have conditions such as depression or anxiety disorders before they present symptoms or seek treatment [7]. Various methods can be employed for screening, including self-reported questionnaires, clinical interviews with mental health professionals, and observing behaviors and physical symptoms. Additionally, studies in affective computing, a field that explores the interaction between human emotions and computational systems, have demonstrated the potential of artificial intelligence (AI) in screening mental disorders [8]. AI-driven technologies can enhance screening by analyzing data from multiple sources, such as text inputs, voice recordings, or facial expressions [9], which is the very focus of this work.
@@
In line with the state-of-the-art, deep models are employed in facial analysis. Pre-trained models are leveraged using two distinct approaches to address the challenge of limited training data (most participants contributed with only a single photo). In a more traditional approach, we employ a transfer learning strategy where convolutional neural networks (CNNs) originally trained for facial expression recognition are fine-tuned for depression-anxiety detection. In a more innovative approach, we propose using powerful vision-language models (VLMs) as a facial analyzer. While VLMs are not suitable for directly assessing depression-anxiety, as demonstrated in our experiments, they excel in zero-shot detailed descriptions of facial expressions that correspond to basic emotions (e.g., anger, happiness, and sadness). This approach detects depression-anxiety by classifying the generated text with simple neural networks instead of directly classifying the image. Experiments conducted under a rigid Leave One Subject Out protocol revealed that our VLM-based approach outperformed the traditional CNNs, achieving an accuracy of 77.6% (+10.0 p.p.) and an F1-score of 56.0% (+11 p.p.).
@@
The rest of this paper is organized as follows. The next section discusses the related works. Section 3 describes the two approaches for AI-driven screening methodology. Section 4 presents the experimental setup, while Section 5 discusses the results. Finally, Section 6 concludes the paper and discusses future works opportunities.
@@ 2 Related Work
The use of facial features to identify non-basic emotions (depression, stress, engagement, shame, guilt, envy, among others) is becoming increasingly popular in various applications. Nepal et al. [10] highlight several elements explored in this type of analysis, such as facial expression itself, gaze direction, as well as general image characteristics like luminosity and background settings. In this domain, machine learning (ML) plays an essential role, particularly with the use of deep learning models and related techniques, such as transfer learning and attention mechanisms [8]. Regarding models, convolutional neural networks (CNNs) are the most popular choice for image-based analysis of non-basic emotions. For instance, Gupta et al. [11] used CNNs to quantify student engagement in online learning, while Zhou et al. [12] applied them for detecting depression based on facial images. Transfer learning plays a crucial role as pre-trained models for tasks like basic facial expression recognition can be fine-tuned for specific targets like detecting stress [13]. Usually, pre-training leverages larger datasets, which is particularly useful when the target dataset is small, as is often the case in mental health applications. Attention mechanisms are also relevant in this context, as they can help models focus on specific regions of the face that are more informative for the target task [14, 15].
@@
Despite the promising results, most studies on depression detection through facial expressions involve capturing images in controlled environments where individuals follow a predetermined script, which limits the realism of the facial expressions [16, 17] and the broad applicability of pre-trained models. Nonetheless, a few studies focus on natural images captured by smartphone cameras for ubiquitous screening. Darvariu et al. [18] developed an application where users record their emotional state through photographs taken by the rear camera of smartphones. Alternatively, front-facing cameras can facilitate the capture of facial images, whose analysis can offer valuable insights into a person’s emotional state [19]. In a recent work, Nepal et al. [10] introduced a depression screening approach named MoodCapture, which was evaluated on a large dataset of facial images collected in the wild (over 125,000 photos). A key aspect of MoodCapture is the passive collection of images in multiple shots while users complete a self-assessment depression questionnaire on their smartphones. The authors argue that these images are preferable to traditional selfies as they capture more authentic and unguarded facial expressions.
@@
Similarly to MoodCapture, the proposed work focuses on analyzing selfies captured by smartphone cameras for anxiety and depression screening. However, our study differs from MoodCapture in two major aspects. First, our work targets a specific population: high-risk pregnant patients accompanied by a team of healthcare professionals. This imposes a limitation on the dataset size when compared to crowd-sourced data. Second, our methodology focuses on face-centric selfies rather than analyzing general image aspects. To the best of our knowledge, this is the first study to explore the use of face-centric selfies for anxiety and depression screening in high-risk pregnant patients. Another relevant aspect of our work is the use of vision-language models (VLMs). Unlike traditional models that predict a limited set of basic emotions, VLMs can capture more nuanced emotions such as awe, shame, and emotional suppression [20]. Despite being used for general emotion analysis, the particular application of VLMs for mental health screening is another contribution of our work.
@@ 3 AI-Driven Screening via Selfie Analysis
This study relies on selfies taken by pregnant patients paired with their responses to the Patient Health Questionnaire-4 (PHQ-4), a four-item instrument for brief screening of anxiety and depression. The PHQ-4 was chosen for this study due to its efficiency and brevity in screening for both anxiety and depression, making it particularly useful in clinical settings where time is limited. Comprising only four items (Figure 1), it integrates questions from the PHQ-2 and GAD-2, which are validated tools for detecting depression and anxiety, respectively. This dual focus allows for a rapid assessment of two of the most prevalent mental health conditions, while maintaining strong psychometric properties, making it a reliable and practical tool for screening purposes in a clinical sample. The total score, calculated as the sum of individual scores, ranges from 0 to 12, with higher scores indicating greater severity of anxiety and depression symptoms. A score of 6 or higher on the PHQ-4 is typically used as a cut-off point for identifying cases where either anxiety or depression (or both) may be present and warrant further clinical evaluation [21]. This cut-off point helps to identify individuals with moderate to severe symptoms of either condition, ensuring efficient screening, including mixed anxiety-depression states, which are fairly common among pregnant women. It allows for initial assessment without the need to differentiate between the two disorders [22].
To enable AI-driven screening, we propose a methodology that leverages selfies and PHQ-4 responses to train machine learning (ML) models for depression-anxiety detection. Figure 2 provides a joint overview of the VLM- and CNN-based approaches addressed in this work. There are two pipelines: the training pipeline (top flow) and the test pipeline (bottom flow). The training pipeline leverages image and PHQ-4 data to train an ML model for depression-anxiety detection. The training requires an image dataset comprising faces and the respective labels (0-normal, 1-abnormal). To assemble this dataset, the first step is manually filtering out invalid selfies, which are those taken with the rear-facing camera or with faces covered by a mask. The face region of the remaining samples is subsequently cropped by using a multi-task cascaded convolutional network (MTCNN) [23]. Selfies with multiple detected faces are discarded, as the PHQ-4 responses are individual. As previously explained, a label is derived for each face image by thresholding the PHQ-4 overall score.
@@ 3.1 CNN-based Approach
This approach employs CNNs for predicting depression-anxiety directly from image data, as illustrated in Figure 2. Training from scratch is unfeasible in our context due to the reduced number of selfies/PHQ-4 responses (147 samples). To circumvent this issue, we start with CNN models pre-trained on the ImageNet dataset and fine-tune them on large-scale facial expression recognition (FER) datasets, which are closely related to our target task. A second fine-tuning is performed to adapt the FER pre-trained models to our task.
This work investigates the adaptation of models pre-trained on the FER2013 [24], and RAF-DB [25] datasets, both of which encompass seven basic emotions (classes): anger, disgust, fear, happiness, sadness, surprise, and neutral state. Four CNN architectures were investigated: EfficientNetV2 [26], ResNet-18 and ResNet-50 [27], and VGG11 [28]. In the second fine-tuning stage, the classifier consists of a CNN backbone pre-trained on FER or RAF-DB appended with a 2-output fully connected layer (classification head). The backbone is frozen during training to prevent overfitting, ensuring that only the classification head is trained.
@@ 3.2 VLM-based Approach
This approach leverages large generative vision-language models (VLMs) for facial analysis. Modern VLMs [29] combine visual encoders with large language models (LLMs) and can simultaneously learn from images and text. This enables them to perform various tasks, such as answering visual questions and captioning images. In particular, we benefit from the VLM zero-shot instruction-following ability to produce high-quality descriptions when provided with an image and text prompt. Three modern VLMs were investigated in this work: LLaVA-NeXT [30], which is an improvement on LLaVA [31], Kosmos-2 [32], and the proprietary GPT-4o.
@@
To build the classification model, features are extracted from the text descriptions by using a pre-trained Sentence-BERT [33] model called all-MiniLM-L6-v2, which is designed to embed sentences and paragraphs into a 384-dimensional representation. Two FFNN architectures (default and alternative) are investigated in this work, as expressed in the Equation (1):
@@ 4.1 Data Collection
Selfies and PHQ-4 responses constitute a subset of data collected for a research project (local ethics board consent: CAAE 64158717.9.0000.0065) involving high-risk pregnant patients from the Clinical Hospital of the University of São Paulo, a public university hospital in São Paulo, Brazil. The research focused on pregnant women aged 18 and above who could provide informed consent, had at least an elementary education, and owned a smartphone. Data collection utilized smartphones, with the participants capturing natural photos (selfies) in such a way that their faces were visible. They were responsible for uploading image data to the REDCap platform [34] instance hosted at the university, which was also utilized for completing the PHQ-4 questionnaire. The bi-weekly data collection was structured into three periods beginning at $t_{0}$, $t_{1}$, and $t_{2}$, with a final date of $t_{k}+12$ weeks, where $k=0,1,2$.
Figure 4 shows the data distribution after manually removing invalid samples. A total of 147 samples from 108 subjects are represented in the left chart. Although the same subject was expected to contribute with multiple samples, most subjects contributed with a single sample, while one contributed with nine. The right-side chart highlights the imbalance in the data, with a bias towards negative samples (PHQ-4  $<$  6): 106 negative versus 41 positive samples.

### v2

# AI-Driven Early Mental Health Screening: Analyzing Selfies of Pregnant Women

@@ Abstract
Major Depressive Disorder and anxiety disorders affect millions globally, contributing significantly to the burden of mental health issues. Early screening is crucial for effective intervention, as timely identification of mental health issues can significantly improve treatment outcomes. Artificial intelligence (AI) can be valuable for improving the screening of mental disorders, enabling early intervention and better treatment outcomes. AI-driven screening can leverage the analysis of multiple data sources, including facial features in digital images. However, existing methods often rely on controlled environments or specialized equipment, limiting their broad applicability. This study explores the potential of AI models for ubiquitous depression-anxiety screening given face-centric selfies. The investigation focuses on high-risk pregnant patients, a population that is particularly vulnerable to mental health issues. To cope with limited training data resulting from our clinical setup, pre-trained models were utilized in two different approaches: fine-tuning convolutional neural networks (CNNs) originally designed for facial expression recognition and employing vision-language models (VLMs) for zero-shot analysis of facial expressions. Experimental results indicate that the proposed VLM-based method significantly outperforms CNNs, achieving an accuracy of 77.6%. Although there is significant room for improvement, the results suggest that VLMs can be a promising approach for mental health screening.
@@ 1 Introduction
Major depressive disorder (depression) is a prevalent mental health disorder affecting over 280 million people globally and a leading cause of disability worldwide. It contributes significantly to the global disease burden (World Health Organization, 2023). Similarly, anxiety disorders, marked by excessive worry and fear, impact millions of lives, often leading to debilitating effects on daily functioning and quality of life. Particularly, there is a recognized vulnerability to mood disorders during pregnancy, with possible worsening of previous emotional conditions and the emergence of new altered mental states that increase the risk of depression and anxiety (Biaggi et al., 2016). This vulnerability is certainly influenced by the significant hormonal changes of pregnancy. However, it is also important to recognize that this is a time of great physical, psycho-emotional, cultural, and social changes, generating psychic stress and increased anxiety.
Several studies have pointed to the association between depression and anxiety during pregnancy with unfavorable obstetric and neonatal outcomes. These disorders can increase the risk of obstetric complications such as cesarean delivery, preeclampsia, preterm birth, low birth weight, small for gestational age newborns, and newborns with low Apgar scores, indicating lower oxygenation at birth (Li et al., 2021; Nasreen et al., 2019; Dowse et al., 2020; Kurki et al., 2000). These findings underline the importance of making a correct and early diagnosis, allowing for appropriate psychiatric and psychological follow-up during pregnancy to improve both maternal and neonatal outcomes.
For early intervention and improved treatment outcomes, screening processes are essential for identifying individuals who may have conditions such as depression or anxiety disorders before they present symptoms or seek treatment (Thombs et al., 2023). Various methods can be employed for screening, including self-reported questionnaires, clinical interviews with mental health professionals, and observing behaviors and physical symptoms. Additionally, studies in affective computing, a field that explores the interaction between human emotions and computational systems, have demonstrated the potential of artificial intelligence (AI) in screening mental disorders (Kumar et al., 2024). AI-driven technologies can enhance screening by analyzing data from multiple sources, such as text inputs, voice recordings, or facial expressions (Liu et al., 2024c), which is the very focus of this work.
@@
In line with the state-of-the-art, deep models are employed in facial analysis. Pre-trained models are leveraged using two distinct approaches to address the challenge of limited training data (most participants contributed with only a single photo). In a more traditional approach, we employ a transfer learning strategy where convolutional neural networks (CNNs) originally trained for facial expression recognition are fine-tuned for depression-anxiety detection. In a more innovative approach, we propose using powerful vision-language models (VLMs) as a facial analyzer. While VLMs are not suitable for directly assessing depression-anxiety, as demonstrated in our experiments, they excel in zero-shot detailed descriptions of facial expressions that correspond to basic emotions (e.g., anger, happiness, and sadness). This approach detects depression-anxiety by classifying the generated text with simple neural networks instead of directly classifying the image. Experiments conducted under a rigid Leave One Subject Out protocol revealed that our VLM-based approach outperformed the traditional CNNs, achieving an accuracy of 77.6%, which represents a gain of approx. 10.0 percentage points (p.p.) compared to the CNNs, and an F1-score of 56.0%, an improvement of approx. 11.0 p.p..
@@
  • Collection of a dataset [^1] Anonymized descriptions produced by the VLMs and corresponding PHQ-4 responses are publicly available at https://bit.ly/3E1EPKw. comprising selfies and PHQ-4 responses from high-risk pregnant patients.
@@
The rest of this paper is organized as follows. The next section discusses the related works. Section 3 describes the two approaches for AI-driven screening methodology. Section 4 presents the experimental setup, while Section 5 discusses the results. Finally, Section 6 concludes the paper and discusses future work opportunities.
@@ 2 Related Work
The use of facial features to identify non-basic emotions (depression, stress, engagement, shame, guilt, envy, among others) is becoming increasingly popular in various applications. Nepal et al. (2024) highlight several elements explored in this type of analysis, such as facial expression itself, gaze direction, as well as general image characteristics like luminosity and background settings. In this domain, machine learning (ML) plays an essential role, particularly with the use of deep learning (DL) models and related techniques, such as transfer learning and attention mechanisms. As highlighted by Kumar et al. (2024), DL models have become the primary choice for detecting non-basic emotions since 2020, surpassing traditional machine learning and image processing algorithms. Regarding models, convolutional neural networks (CNNs) are the most popular choice for image-based analysis of non-basic emotions. For instance, Gupta et al. (2023) used CNNs to quantify student engagement in online learning, while Zhou et al. (2018) applied them for detecting depression based on facial images. Transfer learning plays a crucial role as pre-trained models for tasks like basic facial expression recognition can be fine-tuned for specific targets like detecting stress (Voleti et al., 2024). Usually, pre-training leverages larger datasets, which is particularly useful when the target dataset is small, as is often the case in mental health applications. Attention mechanisms are also relevant in this context, as they can help models focus on specific regions of the face that are more informative for the target task (Viegas et al., 2018; Belharbi et al., 2024).
@@
Despite the promising results, most studies on depression detection through facial expressions involve capturing images in controlled environments where individuals follow a predetermined script, which limits the realism of the facial expressions (Liu et al., 2022; Kong et al., 2022) and the broad applicability of pre-trained models. Nonetheless, a few studies focus on natural images captured by smartphone cameras for ubiquitous screening. Darvariu et al. (2020) developed an application where users record their emotional state through photographs taken by the rear camera of smartphones. Alternatively, front-facing cameras can facilitate the capture of facial images, whose analysis can offer valuable insights into a person’s emotional state (Wang et al., 2015). In recent work, Nepal et al. (2024) introduced a depression screening approach named MoodCapture, which was evaluated on a large dataset of facial images collected in the wild (over 125,000 photos). A key aspect of MoodCapture is the passive collection of images in multiple shots while users complete a self-assessment depression questionnaire on their smartphones. The authors argue that these images are preferable to traditional selfies as they capture more authentic and unguarded facial expressions.
@@
Similarly to MoodCapture, the proposed work focuses on analyzing selfies captured by smartphone cameras for anxiety and depression screening. However, our study differs from MoodCapture in two major aspects. First, our work targets a specific population: high-risk pregnant patients accompanied by a team of healthcare professionals. This imposes a limitation on the dataset size when compared to crowd-sourced data. Second, our methodology focuses on face-centric selfies rather than analyzing general image aspects. To the best of our knowledge, this is the first study to explore the use of face-centric selfies for anxiety and depression screening in high-risk pregnant patients. Another relevant aspect of our work is the use of vision-language models (VLMs). Unlike traditional models that predict a limited set of basic emotions, VLMs can capture more nuanced emotions such as awe, shame, and emotional suppression (Bian et al., 2024). Despite being used for general emotion analysis, the particular application of VLMs for mental health screening is another contribution of our work.
@@ 3 AI-Driven Screening via Selfie Analysis
This study relies on selfies taken by pregnant patients paired with their responses to the Patient Health Questionnaire-4 (PHQ-4), a four-item instrument for brief screening of anxiety and depression. The PHQ-4 was chosen for this study due to its efficiency and brevity in screening for both anxiety and depression, making it particularly useful in clinical settings where time is limited. Comprising only four items (Figure 1), it integrates questions from the PHQ-2 and GAD-2, which are validated tools for detecting depression and anxiety, respectively. This dual focus allows for a rapid assessment of two of the most prevalent mental health conditions, while maintaining strong psychometric properties, making it a reliable and practical tool for screening purposes in a clinical sample. The total score, calculated as the sum of individual scores, ranges from 0 to 12, with higher scores indicating greater severity of anxiety and depression symptoms. A score of 6 or higher on the PHQ-4 is typically used as a cut-off point for identifying cases where either anxiety or depression (or both) may be present and warrant further clinical evaluation (Caro-Fuentes and Sanabria-Mazo, 2024). This cut-off point helps to identify individuals with moderate to severe symptoms of either condition, ensuring efficient screening, including mixed anxiety-depression states, which are fairly common among pregnant women. It allows for initial assessment without the need to differentiate between the two disorders (Javadekar et al., 2023).
To enable AI-driven screening, we propose a methodology that leverages selfies and PHQ-4 responses to train machine learning (ML) models for depression-anxiety detection. Figure 2 provides a joint overview of the VLM- and CNN-based approaches addressed in this work. There are two pipelines: the training pipeline (top flow) and the test pipeline (bottom flow). The training pipeline leverages image and PHQ-4 data to train an ML model for depression-anxiety detection. The training requires an image dataset comprising faces and the respective labels (0-normal, 1-abnormal). To assemble this dataset, the first step is manually filtering out invalid selfies, which are those taken with the rear-facing camera or with faces covered by a mask. The face region of the remaining samples is subsequently cropped by using a multi-task cascaded convolutional network (MTCNN) (Zhang et al., 2016). Selfies with multiple detected faces are discarded, as the PHQ-4 responses are individual. As previously explained, a label is derived for each face image by thresholding the PHQ-4 overall score.
@@ 3.1 CNN-based Approach
This approach employs CNNs for predicting depression-anxiety directly from image data, as illustrated in Figure 2. Training from scratch is unfeasible in our context due to the reduced number of selfies/PHQ-4 responses: 147 samples from a total of 108 participants. To circumvent this issue, we start with CNN models pre-trained on the ImageNet dataset and fine-tune them on large-scale facial expression recognition (FER) datasets, which are closely related to our target task. A second fine-tuning is performed to adapt the FER pre-trained models to our task.
This work investigates the adaptation of models pre-trained on the FER2013 (Goodfellow et al., 2013), and RAF-DB (Li et al., 2017) datasets, both of which encompass seven basic emotions (classes): anger, disgust, fear, happiness, sadness, surprise, and neutral state. Four CNN architectures were investigated: EfficientNetV2 (Tan and Le, 2021), ResNet-18 and ResNet-50 (He et al., 2016), and VGG11 (Simonyan and Zisserman, 2014). In the second fine-tuning stage, the classifier consists of a CNN backbone pre-trained on FER or RAF-DB appended with a 2-output fully connected layer (classification head). The backbone is frozen during training to prevent overfitting, ensuring that only the classification head is trained.
@@ 3.2 VLM-based Approach
This approach leverages large generative vision-language models (VLMs) for facial analysis. Modern VLMs (Bordes et al., 2024) combine visual encoders with large language models (LLMs) and can simultaneously learn from images and text. This enables them to perform various tasks, such as answering visual questions and captioning images. In particular, we benefit from the VLM zero-shot instruction-following ability to produce high-quality descriptions when provided with an image and text prompt. Three modern VLMs were investigated in this work: LLaVA-NeXT (Liu et al., 2024a), which is an improvement on LLaVA (Liu et al., 2024b), Kosmos-2 (Peng et al., 2023), and the proprietary GPT-4o.
@@
To build the classification model, features are extracted from the text descriptions by using a pre-trained Sentence-BERT (Reimers and Gurevych, 2019) model called all-MiniLM-L6-v2, which is designed to embed sentences and paragraphs into a 384-dimensional representation. Two FFNN architectures (default and alternative) are investigated in this work, as expressed in the Equation (1):
@@ 4.1 Data Collection
Selfies and PHQ-4 responses constitute a subset of data collected for a research project (local ethics board consent: CAAE 64158717.9.0000.0065) involving high-risk pregnant patients from the Clinical Hospital of the University of São Paulo (São Paulo, Brazil). The research focused on pregnant women aged 18 and above who could provide informed consent, had at least an elementary education, and owned a smartphone. Data collection utilized smartphones, with the participants capturing natural photos (selfies) in such a way that their faces were visible. They were responsible for uploading image data to the REDCap platform (Harris et al., 2009) instance hosted at the University of São Paulo, which was also utilized for completing the PHQ-4 questionnaire.
For this study, 108 participants were recruited and instructed to submit bi-weekly reports, each consisting of the PHQ-4 questionnaire [^2] The Portuguese version of the PHQ-4 used in this work is provided by Pfizer (Copyright©, 2005 Pfizer Inc). and a selfie. Figure 4 illustrates the data distribution after manually removing invalid selfies. The left chart displays a total of 147 selfies from the 108 participants. Although each participant was expected to contribute multiple samples (11 in total), most participants submitted only a single selfie. The right-side chart highlights the imbalance in the data, with a bias towards negative samples (PHQ-4  $<$  6): 106 negative versus 41 positive samples.